OpenStudy (anonymous):

A bag contains 1 blue, 2 white, and 2 red marbles. Match the columns. Drag the letter after a compound event in the left column and drop it before the matching probability in the right column. Select Submit to see if you are correct. 1. the probability of red and red, with replacement 2. the probability of white and white, without replacement 3. the probability of blue and red, with replacement 4. , without replacement a.16% b.8% c.10% d.20%

OpenStudy (anonymous):

@Directrix plz help

OpenStudy (anonymous):

pretty plz

Directrix (directrix):

Have you worked on this?

Directrix (directrix):

@_Bugzbunny_

OpenStudy (anonymous):

yeah i am just confused

Directrix (directrix):

Prob (RR) with replacement = 2/5 * ?

OpenStudy (anonymous):

2/5 i think because red to red

Directrix (directrix):

2/5 * 2/5 = ?

OpenStudy (anonymous):

so 1- 16%

Directrix (directrix):

4/25

Directrix (directrix):

Why are you subtracting from 1

OpenStudy (anonymous):

i am not the first one is 16%

OpenStudy (anonymous):

am it right

OpenStudy (anonymous):

i not it lol

OpenStudy (anonymous):

8%

OpenStudy (anonymous):

right?

Directrix (directrix):

Yes.

Directrix (directrix):

the probability of blue and red, with replacement 2/5 * 1/5 =

OpenStudy (anonymous):

also 8 % percent but u cant do that

Directrix (directrix):

I don't know what you mean by that.

OpenStudy (anonymous):

i dont think both of them could be the same answer

OpenStudy (anonymous):

of 8%

Directrix (directrix):

B, R no replace 1/5 * 2/4 =

OpenStudy (anonymous):

10%

Directrix (directrix):

This is the one that is wrong. the probability of white and white, without replacement 2 white 5 marbles 2/5 * 1/5 = should be 2/5 * 1/4 = ?

OpenStudy (anonymous):

omg its all in order lol

Directrix (directrix):

That looks like another 10%, so I don't about that, B. Bunny

Directrix (directrix):

Is that it?

OpenStudy (anonymous):

so r they both 10%

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

1is 16% 2 is 10% and is 3 also 10%

Directrix (directrix):

1. the probability of red and red, with replacement 2. the probability of white and white, without replacement 3. the probability of blue and red, with replacement 4. , without replacement

Directrix (directrix):

I do not like percentages. They are dead ends. I prefer the fractional probabilities. They do not mask from whence they came.

OpenStudy (anonymous):

what do u mean

Directrix (directrix):

What is the answer on 4

OpenStudy (anonymous):

i think maybe 20 percent because thats all we have left

OpenStudy (anonymous):

do i just do 1 is a 2 is b 3 is c 4 is d right???

OpenStudy (anonymous):

my brother is being a brat!!! right now and it is not helping my day get better

Directrix (directrix):

Check those.

OpenStudy (anonymous):

1. 16 % 2. 8% 3. 8% 4. 10%

Directrix (directrix):

I wish you'd help me check the probabilities.

OpenStudy (anonymous):

i believe i am correct and the calculator says so

OpenStudy (anonymous):

my teacher says i can use a calculator

Directrix (directrix):

1) 4/25 2) 2/20 3) 2/25 4) 2/20

OpenStudy (anonymous):

so 2 and 4 are 10 %

Directrix (directrix):

GIGO on those calculators

OpenStudy (anonymous):

what

OpenStudy (anonymous):

what is GIGO

Directrix (directrix):

I think the last set of probs are correct, Bunny.

Directrix (directrix):

GIGO http://en.wikipedia.org/wiki/Garbage_in,_garbage_out

OpenStudy (anonymous):

thank u for the medal

Directrix (directrix):

You don't need a calculator to do any of the percentages. Because all the denominators are factors of 100

Directrix (directrix):

You are welcome.

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

got a 98% on the quiz that last one we answered was wrong

OpenStudy (anonymous):

still good

OpenStudy (anonymous):

tome for cornell notes and another quiz the another :(

OpenStudy (anonymous):

time i mean

Directrix (directrix):

I don't know what Cornell Notes are.